RC Update and Run Plan for Aug 20, 2020
RC Meeting Information
RC Daily Meetings are at 4:15 PM, remote
- Todd's Access, He3 refill at 6:45pm
- Reference cell run: 6 hours
- Magnet polarity change
- Positron run (SHMS at 11 degree, P=+7.5 GeV): 1.5 hours
- Moller is still going
Pending problem(s) / Things to track
Recent polarization measurements
- OWL (08/20) 2:30 - 29%
- OWL (08/20) 12:20 - 10%
- OWL (08/20) 14:51 - 19.5%
Current Run Plan
Normal Production running
- Make sure the raster size is correct (4--4.5mm)
- Monitor beam parameters/configuration very often and post an entry like this at the beginning of your shift and when run conditions have changed
- Running on cell targets: raster always ON (4--4.5mm, size will be determined by the scraping checks), orbit locks ON, current ramp rate 1 muA/sec
- Always take 120Hz of EDTM
- Fill the BTA with realistic numbers. If the desired beam current is not delivered prorate your ABUs!
- Change the IHWP at least every 8 hours when running on glass cells. If Hall A hasn't changed in 8 hours, call the RC
- Run plan
- Complete the Moller measurement (still 4-6 hours of work remain), while keeping an eye on the target polarization.
Notices in Effect for Shift Crew
- Follow all COVID rules and regulations as well as remind each other.
- Beam positions: 
- Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
- Inform the RC of any access that the Target Team requires. The RC will coordinate with RadCon to make everything run smoothly.
- Please make sure that the orbit locks are enabled and the beam does not drift from its nominal position.
- Ramp rate lock instructions are here and added to checklist
- An new reference Cell refill procedure: New procedure
- Magnet policies
- Spectrometer GUI (
go_magnets) should always be visible on Shift Leader computer, hcdesk3
- Slow ramping procedure of the SHMS, go small steps to the MOL, an great example (1 hour 30 min process)
- HMS Q1 must be ramped down in 50A steps in all cases
- Do not ramp HMS Q2 to zero, only go to 10A and then switch off power.
- If any of the Spectrometers magnets trip, take a screenshot of the PSU window and check the hall for smoke or water with the camera before resetting the interlock and post it to hclog.
- Spectrometer GUI (
- New SHMS matrix element
Planned Accelerator Configuration Changes
- Hall C to 1st pass on 8/31 and back to 5th pass on 9/03
- Bluejeans remote connection information:
To join the Meeting: https://bluejeans.com/861439920 To join via Room System: Video Conferencing System: bjn.vc -or-184.108.40.206 Meeting ID : 861439920 To join via phone : 1) Dial: +1.408.740.7256 (US (San Jose)) +1.888.240.2560 (US Toll Free)) +1.408.317.9253 (US (Primary, San Jose)) (see all numbers - http://bluejeans.com/numbers) 2) Enter Conference ID : 861439920